OpenStudy (anonymous):

Solve. 3/x=1/x-2 a. x=3 b.x=4 c. x=6

Parth (parthkohli):

Eliminate fractions. Do you know how to do it?

OpenStudy (anonymous):

nah

Parth (parthkohli):

Okay, if: \(\Large \color{midnightblue}{\rightarrow {a \over b} = {x \over y}, then \rightarrow ay = bx }\)

Parth (parthkohli):

x = 3(x - 2) by cross multiplying x = 3x - 6 6 = 2x x = 3

OpenStudy (anonymous):

thank u

Parth (parthkohli):

YW, YG :P